OpenStudy (anonymous):

I need help working this out! Question: Solve the following system of equations. x+y-4z=-7 2x+3y-3z=-6 2x+2y-z=0 I know that my first step is adding two equations together to get x by it's self. So I took 2x+3y-3z=-6 and multiplied it by -1. So that I could add it to 2x+2y-z=0, and get the x by it's self. Now if that was right I got x-1y+2z=6 as the equation and I need help as to what to do next and if that was right, I'm pretty sure next I add that equation with the first? Please help! Thanks!

OpenStudy (anonymous):

I like what you did about multiplying the equation by -1 so you can add it (you could have just subtracted) but you need to re-add them. The x's should disappear

OpenStudy (anonymous):

What do I re-ad?

OpenStudy (anonymous):

You want to add 1 and 3 also 2 and 3 together. The goal is to eliminate a variable so that you are left with 2 equations and 2 variables. Re-add the equations. You need to start over

OpenStudy (anonymous):

okay thank so i should start by adding 1 and 3?

OpenStudy (anonymous):

Step 1: Get a variable to have all the same coefficients ( x+y-4z=-7 )*2 2x+3y-3z=-6 2x+2y-z=0

OpenStudy (anonymous):

Step 2: Eliminate that variable by subtacting (or adding the opposite if that is easier for you) I am going to subtract 3 from 1 and 2 from 1 1) 2x+2y-8z=-14 2) 2x+3y-3z=-6 3) 2x+2y-z=0

OpenStudy (anonymous):

okay so i should multiply the first by 2 alright ill do that all now thanks so much

OpenStudy (anonymous):

I's so confused right now how do i get the variable by it's self

OpenStudy (anonymous):

1) 2x+2y-8z=-14 3) 2x+2y-z=0 a)-7z=-14 1) 2x+2y-8z=-14 2) 2x+3y-3z=-6 b) -y-5z=-8 Step 3: Now that we have 2 equations and 2 unknowns we can solve for those unknowns. I am going to use substitution a) -7z=-14 b) -y-5z=-8 a) z=2 b) -y-5(2)=-8 y=-2

OpenStudy (anonymous):

Oh! and x would be 3!

OpenStudy (anonymous):

after you sub the y into the first equation, right?

OpenStudy (anonymous):

Step 4: Using the values we just found we need to plug them into one of the original equations to solve for the last unknown. I always choose the easiest one. 1) x+y-4z=-7 x+(-2)-4(2)=-7 x-2-8=-7 x=3

OpenStudy (anonymous):

Thank you so much! I was stuck on that problem forever and you really helped!

OpenStudy (anonymous):

Sorry I wasn't reading your comments. Tell me what steps confuse you and I'll try to explain

OpenStudy (anonymous):

yw. Do you get it?

OpenStudy (anonymous):

yes after you explained the first two on what y and z was I plugged in y and got that x was 3

OpenStudy (anonymous):

Yup. Good job. Feel free to tag me in future questions by doing "@ChmE"

OpenStudy (anonymous):

I have a small one I'll post in a second.